### v1

# Sound Field Synthesis with Acoustic Waves  [^1] Author is affiliated with Amazon Inc. USA.  [^2] This manuscript is an expanded version of a conference paper of the same title

Abstract
We propose a practical framework to synthesize broadband sound-field on a small rigid surface based on the physics of sound propagation. The sound-field is generated as a composite map of two components: room component and device component; with acoustic plane waves as core element. This decoupling of room and device components significantly reduces the problem complexity and provide accurate rendering of the sound-field. We describe in details the theoretical foundations, efficient procedures, and engineering applications of the proposed framework. The effectiveness of the proposed framework is established through rigorous validation under different environment setups.
Index Terms:
Room Acoustics, Sparse Recovery, Acoustic Wave Decomposition.
I Introduction
Sound-field synthesis at a microphone array in a room is the process of synthesizing audio at each microphone of the array from a source signal emanating from a sound source elsewhere in the room. It is a key task in evaluating performance metrics of speech/audio communication devices, as it is a cost-effective methodology for data generation to replace real data collection, which is usually a slow, expensive, and error-prone procedure. Acoustic modeling techniques are frequently utilized to generate synthetic data to either replace or augment real data collection at a fraction of data collection cost. These techniques usually aim at estimating the Room Impulse Response (RIR) between two points in a room. The RIR is either computed empirically using direct measurement, or simulated using a model for room acoustics. Empirical methods are in general accurate, but they are relatively expensive because of the required human labor. The whole procedure for RIR calculation has to be repeated for every source/receiver position combination and for every new device with a different form factor. Moreover, empirical methods require the hardware availability of the device under test which is not always possible especially during hardware development.
Simulation methods address the above issues by relying on computational acoustics rather than measurements to estimate the RIR. A brute-force simulation would solve the inhomogeneous acoustic wave equation with proper boundary conditions of the room and device surface [1]. Though theoretically viable, it requires significant effort to characterize all boundary conditions in a typical room. Further, the simulation time can be prohibitive if it is evaluated over a broadband spectrum. Moreover, the whole simulation needs to be repeated for every new form factor of the device under test. To address the computational complexity, the image source method [2] has been widely used to approximate point-to-point room acoustics. It utilizes the ray tracing concept [1] to significantly reduce modeling and computational complexity of brute-force simulation. Though simple and effective in some scenarios, the image source method has few limitations. For example, it has poor approximation at low frequencies, and it cannot model small surfaces, e.g., furniture, and rough surfaces, e.g., curtains.
In this work, we describe a novel procedure that combines empirical and simulation methods to provide a balanced tradeoff between the two approaches for sound-field synthesis. It splits the sound-field into two independent components: room component, and device component; and the overall sound-field is the composite mapping of the two components. The room component abstracts the room impact at an interior point, due to a predefined sound source in the room, as a superposition of acoustic plane waves; which is computed using a single measurement with a large microphone array. The device component is computed using acoustic simulation or anechoic measurements to evaluate the fingerprint of each acoustic plane wave on the device surface (as measured at the microphone array of the device). The overall acoustic pressure on the device surface when placed at an interior point in a room is computed by plugging in the computed device fingerprints into the acoustic plane wave representation at that point. This arrangement abstracts room acoustics and allows reusing room information with different devices when tested in the same room. Therefore, it enables the concept of room database, which contains abstract room acoustics information that is independent of the device under test. Likewise, it allows reusing the same device component with different rooms. To enable the proposed method, we develop a general procedure to compute the plane wave decomposition at a point in a room by applying sparse recovery techniques on an audio capture with a large microphone array. We also introduce the device dictionary concept, that captures the device acoustic behavior; which is a general concept for general microphone array processing systems [3]. The proposed methodology is rigorously validated across many rooms and many devices with different form factors and microphone array configuration. The synthetic RIR is shown to match the true RIR, in the least square sense, over a broadband spectrum up to $8$ kHz. The synthesis methodology is also shown to closely resemble real measurements in evaluating higher level metrics, e.g., word error rate, and false rejection rate.
The acoustic plane wave expansion has been used in earlier work with model-based sound-field reconstruction, e.g., [4, 5, 6, 7]; where acoustic plane waves are used as kernels for sound-field reconstruction. The plane-wave expansion is interpolated with free-field propagation model to reproduce the sound field within a convex source-free zone. The plane-wave expansion is computed from measurements of an array of microphones placed at the zone perimeter. The computation of the expansion is done either through spherical harmonics or using sparse recovery techniques. In this work, we study a different problem of reproducing the sound field on a rigid surface that is placed at the same point in the room. The key contribution of the current work is introducing the device dictionary concept; which captures the impact of acoustic plane waves on the device surface. This enables the generalization of the sound-field production to a rigid surface with an arbitrary form factor.
The paper is organized as follows. In section II, we lay down the theoretical foundation of the work. The details of the proposed framework are described in section IV. Then, we present the validation results in section V. Finally, we describe few engineering applications in section VI. The following notations are used throughout the paper. A bold lower-case letter denotes a column vector, while a bold upper-case letter denotes a matrix. $M$ always refers to the number of microphones. The independent variables $t$ and $\omega$ refer to time and frequency respectively. Additional notations are introduced when needed.
II Foundations
II-A Acoustic Plane Waves
Acoustic plane waves are eigenfunctions to the homogenous Helmholtz equation, and they constitute a powerful tool for analyzing the wave equation. Further, plane waves are a good approximation of the wave-field emanating from a far-field point source [8]. The acoustic pressure of a plane-wave with vector wave number $\bf{k}$ is defined at a point ${\bf{r}}=(x,y,z)$ in the three dimensional space as:
$\psi({\bf{k}})\triangleq p_{0}e^{-j{\bf{k}}^{T}{\bf{r}}}$ (1)
where ${\bf{k}}$ is the three-dimensional wavenumber vector.
The plane-wave decomposition (PWD) has been used for approximating point-source seismic recording [9, 10, 11], and sound field reproduction [12, 13, 14, 8]. A local solution to the homogenous Helmholtz equation can be approximated by a linear superposition of plane waves of different angles of the form [15, 10]:
$p(\omega)=\sum_{l\in\Lambda}\alpha_{l}\ \psi\left({\bf{k}}(\omega,\theta_{l},\phi_{l})\right)$ (2)
where $\Lambda$ is a set of indices that defines the directions of plane waves $\{\theta_{l},\phi_{l}\}$, each $\psi({\bf{k}})$ is a plane wave as in (1), and $\{\alpha_{l}\}$ are complex scaling factors. We will refer to the wave-field in (2) as the overall background acoustic pressure. The decision variables in this approximation are $\left\{\Lambda,\{\alpha_{l}\}_{l\in\Lambda}\right\}$.
III Device Acoustic Dictionary
Generalizing the free-field plane-wave expansion in (2) to include the scattering due to the device surface, requires computing the device acoustic response to each plane wave. The device response to all plane-wave in the three-dimensional space is collectively referred to as the device acoustic dictionary. The total wave-field at any point on the device surface when the device is impinged by an incident plane-wave $p_{i}({\bf{k}})$ has the general form:
$p_{t}=p_{i}+p_{s}$ (3)
where $p_{t}$ and $p_{s}$ refer to the total and scattered wave-field respectively. $p_{t}$ can be computed numerically by inserting (3) in the Helmholtz equation and solving for $p_{s}$ with appropriate boundary conditions. If a microphone array of size $M$ is mounted on the device surface, and the microphone port size is much smaller than the wavelength, then each microphone can be approximated by a point on the device surface. In this case, the total field, ${\mathbf{p}}_{t}$, at the microphone array, due to an incident plane wave $p_{i}({\mathbf{k}})$, is a vector of size $M$ whose entries are the corresponding total field at each individual microphone. The device acoustic dictionary of a device is composed of vectors of total acoustic wave-field.
An entry of the device dictionary can be either measured in anechoic room with single-frequency far-field sources, or computed numerically by solving the Helmholtz equation on the device surface with background plane-wave using the device CAD model. Both methods yield same result, but the numerical method has much lower cost and it is less error-prone because it does not require human labor.
For the numerical method, each entry in the device dictionary is computed by solving the Helmholtz equation, using Finite Element Method (FEM) or Boundary Element Method (BEM) techniques, for the total field at the microphones with the given background plane wave. The device CAD is used to specify the surface, which is modeled as sound hard boundary. To have a true background plane-wave, the external boundary should be open and non-reflecting. In our model, the device is enclosed by a closed boundary, e.g., a cylinder or a spherical surface. To mimic open-ended boundary we use Perfectly Matched Layer (PML), which defines a special absorbing domain that eliminates reflection and refractions in the internal domain that encloses the device [16]. Standard packages for solving partial differential equations, e.g., [17] are used, and the simulation is rigorously validated with measured acoustic pressure on different form-factors. In Fig. 1, we show an example of the frequency amplitude and phase of the inter-channel transfer function of both simulated and anechoic measured response for a microphone array mounted on a sphere. The reference channel in the inter-channel transfer function is the first microphone that is hit by the plane wave. The phase plot at the bottom is the phase error between the measured and simulated response. In the ideal case, the magnitude response of the measured and simulated transfer function should coincide, and the phase error is identically zero. The matching of the magnitude response is quite clear and ripples in the measure response is due to the impact of minor reflections in the anechoic room. Similarly, the phase error cannot be identically zero in practice because of the finite geometric precision in the position in the anechoic room, which results in unavoidable linear phase error that is shown in the phase plot. More validation examples were described in [3]. In [18, 19], comparisons between simulated and theoretical acoustic pressure responses were presented.
Figure 1: (top) Measured (solid line) and simulated (dotted line) total field of a microphone array mounted on a sphere of PW, (bottom) phase error, with azimuth = $150^{\circ}$, elevation = $90^{\circ}$.
The output of the above model is the plane-wave dictionary of the device
$\mathcal{D}\triangleq\{\boldsymbol{\beta}_{l}(\omega)\triangleq{\bf{p}}_{t}(\omega,\theta_{l},\phi_{l}):\forall\ \omega,l\}$ (4)
where each entry in the dictionary is a vector of length $M$, and each element in the vector is the total acoustic pressure at one microphone in the microphone array when a plane wave with ${\bf{k}}(\omega,\theta_{l},\phi_{l})$ hits the device. The dictionary also covers all frequencies of interest typically up to $8$ kHz.
IV Proposed Framework
IV-A Overview
The key idea in the proposed sound synthesis methodology is to generalize the plane-wave expansion in (2), which summarizes room acoustics at a point in a room, to include the impact of scattering due to the device surface. This generalization would yield the combined effect of the room and the device under test, which is the final objective. If the device dimensions are much smaller than the room dimensions, then secondary reflections due to the device surface are negligible; and the device impact on the room acoustics could be ignored. Hence, even after introducing the device into the room, the directions and weights of the acoustic plane waves in (2) do not change. However, because of the device surface, each plane wave $\psi\left({\bf{k}}(\omega,\theta_{l},\phi_{l})\right)$ in (2), has an acoustic fingerprint at the device microphone array; which is the corresponding entry in the device dictionary, $\boldsymbol{\beta}_{l}(\omega)$. Hence, if the device is placed at a point in the room whose sound-field is expressed as in (2), then the observed sound field vector at the device microphone array is
${\mathbf{p}}(\omega)=\sum_{l\in\Lambda}\alpha_{l}\ \boldsymbol{\beta}_{l}(\omega)$ (5)
The transition from the sound-field in (2) in the absence of the device to the sound field in (5) in the presence of the device is the technical foundation of the proposed synthesis framework. This transition is enabled by the linearity of the wave equation and the introduction of the device acoustic dictionary as described in section III.
Note that, if another device with acoustic dictionary $\mathcal{D}^{(2)}\triangleq\{\boldsymbol{\beta}_{l}^{(2)}(\omega):\forall\ \omega,l\}$ is placed at the same point, then the observed sound-field at the second microphone array
${\mathbf{p}}^{(2)}(\omega)=\sum_{l\in\Lambda}\alpha_{l}\ \boldsymbol{\beta}_{l}^{(2)}(\omega)$ (6)
where only the mapping through device dictionary changes, while the directions, $\Lambda$, and weights, $\{\alpha_{l}\}$, of constituent plane waves do not change. This is the essence of the proposed methodology that separates room and device components.
Hence, the three steps for sound-field synthesis at a microphone array mounted on a device that is placed at a point in the room are as follows:
  1. 1.
    Compute the free-field plane wave expansion at a point as in (2). This summarizes room acoustics at a point in the room. A procedure for computing this expansion is described in section IV-B. This process is repeated for every new point in a room, and it is independent of the device under test.
  2. 2.
    Generate the acoustic dictionary of the device under test as described in section III. This is computed once per device and it is independent of the room.
  3. 3.
    For each room position, combine the plane-wave expansion with the device dictionary as in (5) to synthesize the sound-field at the device microphone array.
Repeating step $1$ above for multiple rooms and multiple positions within each room generates a room database; which is generated only once, but could be reused in evaluating and generating data for new devices.
IV-B Acoustic Plane Wave Decomposition
The main technical hurdle in the proposed framework is computing the plane-wave expansion (2) at a point in a room with a source signal emanating from another point in the room. In the proposed framework, this is computed via data capture with a large microphone array of $32$ microphones mounted on a sphere (EigenMike) [20]. The large microphone array is necessary to avoid having an underdetermined system of equations in recovering the constituent plane waves. It was found experimentally that $20$ to $30$ plane waves are sufficient for an accurate approximation of the sound field (with reconstruction error less than $-30$ dB). The plane wave decomposition is composed as an optimization problem whose objective is minimizing the difference in the least square sense between observed and synthesized sound fields. If the observed sound field of the EigenMike at frequency $\omega$, is ${\mathbf{y}}(\omega)$, then the objective function to be minimized has the form
$J=\int_{\omega}\|{\mathbf{y}}(\omega)-\sum_{l\in\Lambda}\alpha_{l}(\omega)\bar{\boldsymbol{\beta}}_{l}(\omega)\|^{2}+\lambda\sum_{l\in\Lambda}|\alpha_{l}(\omega)|$ (7)
where $\left\{\bar{\boldsymbol{\beta}}_{k}(\omega)\right\}$ are the entries of the EigenMike acoustic dictionary at frequency $\omega$. The decision variables are the set of indices $\Lambda$, and the corresponding weights $\{\alpha_{l}\}$. The L1-regularization in (7) is added to stimulate a sparse solution as $|\Lambda|<30$ is much smaller than the dictionary size, which is in the order of $10^{3}$. The objective function can be put in matrix form as:
$J=\int_{\omega}\|{\mathbf{y}}(\omega)-{\mathbf{A}}(\omega)\ .\ \boldsymbol{\alpha}\|_{2}^{2}+\lambda\ |\boldsymbol{\alpha}|_{1}.$ (8)
where ${\mathbf{A}}$ is a matrix whose columns are the individual entries of the acoustic dictionary at frequency $\omega$, i.e., $\{\bar{\boldsymbol{\beta}}_{l}(\omega)\}$. The above problem is a form of the well-known LASSO optimization [21] that is encountered in numerous sparse recovery problems in statistics and signal processing. Many efficient solutions have been proposed for this problem under various conditions [22, 23]. The big microphone array size in the EigenMike provides much flexibility in solving (8) because the observation size is bigger than the number of nonzero components in $\boldsymbol{\alpha}$. Note that, the optimization problem in (8) is solved only once for a given source/receiver position in a room, and it is solved offline. Therefore, it does not have constraints on computational complexity, memory, or latency. In our analysis, the orthogonal matching pursuit algorithm [24] was used to recover $\Lambda$ and $\boldsymbol{\alpha}$, though other existing solutions to the sparse recovery problem can be used with this formation. This was generalized for smaller microphone arrays of arbitrary geometry in [25].
For a given source signal, the above procedure is repeated at each frequency $\omega$, and at each time frame to generate a time-frequency map of the active set $\Lambda(t,\omega)$ and the corresponding weights $\boldsymbol{\alpha}(t,\omega)$. To synthesize the sound field for another device with acoustic dictionary $\{{\boldsymbol{\beta}}_{l}(\omega)\}$ at this particular source/receiver position and source signal, the synthesis formula (5) is applied at each time-frequency cell with the corresponding parameters $\Lambda(t,\omega)$ and $\boldsymbol{\alpha}(t,\omega)$. Generating the sound-field for an arbitrary source signal requires the computation of the room impulse response, which is described in the following section.
IV-C Room Impulse Response (RIR) Computation
RIR aims at modeling the acoustic channel between source and receiver as a linear time-invariant system. The RIR combines both room acoustics and scattering due to device surface, and it is computed once for a given device and a given source/receiver positions in a room. It is a multichannel transfer function where the number of channels equals the size of the microphone array. For RIR computation, a special source signal that covers the whole frequency spectrum, e.g., white noise or Golay sequence [26], is utilized. A source signal with short duration of $2$-$3$ seconds is sufficient for RIR computation. For a source signal $x(t,\omega)$, the EigenMike is utilized to generate the time-frequency map of the plane wave decomposition as described in the previous section. For a device under test, this time-frequency map is combined with the device dictionary to generate the multichannel output signal ${\mathbf{y}}(t,\omega)$ as in (5). The transfer function between $x(t,\omega)$ and ${\mathbf{y}}(t,\omega)$ is computed using system identification techniques. For example, by applying Wiener-Hopf equation in the frequency domain [27], we get
$\hat{\mathbf{h}}(\omega)=\frac{\mathbf{S}_{xy}(\omega)}{S_{xx}(\omega)}$ (9)
where $\hat{\mathbf{h}}(\omega)$ is the acoustic transfer function in the frequency domain, and
$S_{xx}(\omega)$ $=$ $\mathbb{E}\left\{x^{*}(t,\omega)\ x(t,\omega)\right\}$ (10)
${\mathbf{S}}_{xy}(\omega)$ $=$ $\mathbb{E}\left\{x^{*}(t,\omega)\ {\mathbf{y}}(t,\omega)\right\}$ (11)
After RIR estimation for a device at a point in a room, the sound field for an arbitrary source signal $u(t,\omega)$ is computed as
${\mathbf{y}}(t,\omega)=\hat{\mathbf{h}}(\omega)\ .\ u(t,\omega)$ (12)
Note that, the RIR computes only the part of the sound field that is correlated with the source signal, and it disregards the background ambient noise and other interferences in the room. Nevertheless, this ambient noise is uncorrelated with the source signal and its time-frequency map, ${\mathbf{b}}(t,\omega)$, is computed once as described in section IV-A, then added to the sound-field for any source signal. In this case, the synthesized sound-field in (12) is modified to
${\mathbf{y}}(t,\omega)=\hat{\mathbf{h}}(\omega)\ .\ u(t,\omega)+{\mathbf{b}}(t,\omega)$ (13)
IV-D Discussion
The proposed method is a combination of measurements (for room component) and simulation (for device component). A single measurement with a large microphone array is required per room position, and this measurement is reused for all devices. The measurement is processed by plane wave decomposition to compute the time-frequency map of the acoustic decomposition that is combined with device dictionary to generate the total sound field. Similarly, the device dictionary is computed once, and it is combined with any room to generate the full sound-field. The computational complexity for computing the broadband device dictionary is small because it is computed in anechoic setup. Further, it is highly parallelizable because the same process is repeated at different frequencies and at different directions for plane wave. The concept of splitting room acoustics and device acoustics significantly reduces the measurement/simulation overhead. In addition to simplifying both room and device modeling, abstracting room acoustics in a single measurement and device acoustics with the device dictionary enables reuse of either components with the other side.
The proposed approach provides an accurate approximation of room acoustics and alleviates the need for full room simulation whose complexity is prohibitive at high frequencies for a regular-size room. Further, this single room measurement eliminates the need to model the room interior surfaces, which can also be an overly time-consuming process. As compared to the image source method, the proposed method addresses all the limitations outlined in section I as follows:
  1. 1.
    The plane-wave expansion model in (5) is valid at all frequencies.
  2. 2.
    The impact of device scattering is incorporated through the device dictionary in sound-field synthesis.
  3. 3.
    The impact of all boundary conditions in the room is inherently included in the plane-wave expansion in (2). It automatically accounts for all surfaces in the room without explicitly modeling them.
Note that, it is possible to combine the device acoustic dictionary with the image source method [2] to further educe complexity at the cost of lower accuracy. In the image source method, the concept of a sound wave is replaced by sound rays; which is a small portion of a spherical wave with vanishing aperture [1]. These sound rays from the image source method can be regarded as a crude approximation of acoustic plane waves in (2); which eliminates the need for room measurements. If these sound rays are combined with device dictionary, then it extends the image source method to account for scattering due to the device surface. However, it still inherits the other gaps of the image source method as previously outlined in section I.
V Experimental Validation
The first experiment aims at validating the plane wave decomposition procedure as described in section IV-B. In Fig. 2, we show the reconstruction error of the EigenMike for two different source signals versus the number of plane wave in the expansion. The reconstruction error (aka, Goodness of Approximation, GoA) is defined as:
$GoA\triangleq\frac{\int_{\omega}\|{\mathbf{y}}(\omega)-\sum_{l\in\Lambda}\alpha_{l}(\omega)\bar{\boldsymbol{\beta}}_{l}(\omega)\|^{2}}{\int_{\omega}\|{\mathbf{y}}(\omega)\|^{2}}$ (14)
where ${\mathbf{y}}(\omega)$ is the observed sound-field at the EigenMike.
Figure 2: EigenMike reconstruction error versus the number of plane waves in PWD
As noted from the figure, a small number of plane waves is sufficient for sound field approximation with error less than $-20$ dB.
In the following set of experiments, the RIR procedure as described in section IV-C is evaluated. The experiments were conducted in three different rooms with furniture that resemble typical bedroom and living rooms. The experiments were conducted in $24$ different positions within the three rooms. The EigenMike was colocated with four devices with different form factors and microphone array size. Two devices have cylindrical form factor, one has cube-like shape, and the fourth is a slated sphere. The $24$ test positions cover different positions in the room: middle of the room, next to a wall, and at a corner. For each position, the origin of the EigenMike and collocated device were aligned properly using laser beams. The measured and synthetic RIR are computed as described in section IV-C. In all cases, there exist strong resemblance between measured and synthetic RIR at all frequencies, and the reconstruction signal-to-noise ratio (SNR) for the devices is between $19$ to $23$ dB. An example of the transfer function and the impulse response of the measured and synthetic RIR is shown respectively in Figures 3 and 4. This is a typical behavior at all positions and devices.
Figure 3: Frequency response of measured and synthetic RIR for a microphone array of size 4 on spherical cap
Figure 4: Impulse response of measured and synthetic for microphone array of size 4 on spherical cap
In the third set of experiments, we study the impact of mismatch between true and synthetic RIR in evaluating high level data metrics. For this test, we evaluated the False Rejection Rate (FRR) of a keyword in the source signal. The true FRR is computed by processing the device signal after the convolution of the source signal and the true RIR. Similarly, the synthetic FRR is from the convolution of the same source signal and synthetic RIR. The relative absolute FRR is defined as
Relative FRR $\text{Relative FRR Error}\triangleq\frac{|\ FRR_{true}-FRR_{synthetic}\ |}{FRR_{true}}$ (15)
The cumulative density function of the relative absolute error (of all $24$ room positions and all devices) is shown in Fig. 5. From the figure, we notice that the relative error between true and synthetic FRR is less than $10\%$ with $95\%$ probability.
Figure 5: Cumulative density function (CDF) of the relative absolute FRR error
VI Relevant Applications
VI-A Data Generation
The primary application of the proposed sound synthesis procedure is generating synthetic data to compute performance analytics for the device under test, e.g., False Rejection Rate (FRR), Word Error Rate (WER), and audio quality metrics. It provides a low-cost high-quality alternative to real data collection; which is an expensive and time-consuming process. Further, the proposed framework provides control of the environment conditions for customized usage scenarios. The synthetic data can also be used to augment real data collection for training deep learning acoustic models, which require large amount of data under diverse acoustic conditions. Note that, the proposed methodology requires only the device CAD (to compute the device dictionary), and does not need the actual device hardware. Therefore, it could be integrated with the hardware design process to evaluate different metrics at early design stages without building hardware prototypes.
VI-B Microphone Array Processing
The traditional model for microphone array processing in the literature utilizes free-field steering vectors that capture only phase differences due to wave propagation and neglect the magnitude component due to scattering with device surface [28, 29]. This limitation results in undesirable effects with beamforming, e.g., spatial aliasing. Rather, the acoustic dictionary, as described in section III, is a generalization of free-field steering vectors that captures device response to acoustic plane wave. The entries of the acoustic dictionary provide more accurate steering vectors because they have both magnitude and phase components. Therefore, beamformer design with steering vectors from acoustic dictionary provides significantly better directivity in real room setting [3]. Further, the plane wave decomposition in (5) incorporates scattering at the device surface; hence, it provides an accurate acoustic map of the device surrounding. It can be regarded as an invertible spatial transformation that maps microphone array observations into spatial components. This enables many applications related to microphone array processing, e.g., sound source localization, sound source separation, and dereverberation.
VII Conclusions
We presented an effective procedure for sound-field synthesis based on acoustic plane waves. The system is composed of three components:
  1. 1.
    Room component: which abstracts the sound field at a point in a room as a superposition of acoustic plane waves as in (2). The complexity of room acoustics is encountered by a single recording with a large microphone array followed by plane wave decomposition to resolve the acoustic scene. Repeating the process once at multiple rooms generates a room database that is reused with any device.
  2. 2.
    Device component: which computes the response of the device surface to acoustic plane waves in the 3D space. The process is highly parallelizable and the response is computed in anechoic environment which significantly reduces the simulation complexity. The process is done once per device, and the resulting dictionary is then combined by the different rooms in the room database.
  3. 3.
    Synthesis component: which synthesizes the sound field on the device surface by combining the constituent plane waves from the first component with the surface response in the second component as in (5) .
The decoupling of room and device components significantly reduces the complexity by enabling reuse through combining the two components. We presented efficient algorithms from numerical analysis and system theory to realize the proposed system. The merits of the proposed framework and comparison with prior art were outlined in section IV-D.
The contributions of this work are summarized as follows:
  1. 1.
    A methodology for decoupling and combining room acoustics and impact of device surface for accurate sound field realization.
  2. 2.
    A novel algorithm to compute plane wave decomposition of a point in a room based on measurement with a large microphone array with sparse recovery techniques.
  3. 3.
    A methodology for characterizing device acoustics based on response to acoustic plane waves.
  4. 4.
    A framework for data generation based on room impulse response that is derived from the proposed sound synthesis procedure.
We also highlighted few relevant applications with high impact that are based on the proposed framework; which are investigated in details in future publications.